OpenStudy (anonymous):

I need help fixing some reflecting questions: This is my teachers feedback: 6. Think about why the equations came out the same in #4. 7. Since your Spacecraft has a specific starting and ending point we wouldn’t be able to have any values past those points. Think about how that would affect your domain and range. 8. Think back to the airplane example in this lesson and how we couldn’t have a negative value for y. Will your lines always go to infinity or will they have stopping points?

OpenStudy (anonymous):

6. Does the point you select matter when your write a point-slope equation? Explain your reasoning using complete sentences. No, because I could have chose any points, and still put them in point slope form. It would have worked the same, just with a different answer, and equation.

OpenStudy (anonymous):

7. 7.Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences. No, because a line is the set of all points. As long as the point is on the line, it will help in building the line up.

OpenStudy (anonymous):

8. Explain, using complete sentences, why it is important to understand any limitations on the domain and range. The domain and range help you properly graph a equation. It is also important to know, because if you are solving a problem in the real world that have to do with distance, length, or volume, then understanding domain and range will help you overcome the limitations when solving the problem.

OpenStudy (anonymous):

I don't understand what I did wrong...

OpenStudy (anonymous):

When he says #4 this is what he is talking about: In question 2, you selected one of two points (Point A or Point B) to be included in your point-slope equation. Write the point-slope form of that equation again, using the other point’s coordinates. y-y1=m(x-x1) y-5=1/2(x-3)

OpenStudy (amistre64):

does a line pass thru all the points that it passes thru?

OpenStudy (anonymous):

Umm...IDK, these are the points Launch Area:___(1, 2)___ Point A: (1,4) Point B: (3,5) Point C: (5,3)

OpenStudy (amistre64):

if you drive from new york to LA, will you pass thru all the cities that you pass thru?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

a line is a collection of points all lined up one after the other along a given slope we can construct the equation of a line by using its slope and any of the points contained by the line; so it doesnt matter if we use one particular point or another

OpenStudy (anonymous):

So that's what I should have put for number 6?

OpenStudy (amistre64):

something along those "lines" yes .... or at least thats what I would have considered

OpenStudy (anonymous):

Okay...I'll put it in my own words... What about number 7?

OpenStudy (amistre64):

the point you choose is jut a starting point that can get you to all the other points on the line; so as long as you are constructing the equation with a point that is already on the line; the equation will give you the same results is all.

OpenStudy (amistre64):

what do domain and range tell us?

OpenStudy (anonymous):

a function is a relation between a set of inputs and a set of permissible outputs with the property that each input is related to exactly one output.

OpenStudy (amistre64):

soo, domain and range deal with a set of inputs and outputs right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

if we restrict the domain and range, we restrict the values that we can use to apply to our setup. Do we want values that go beyond what we are trying to reach? or values that allow us to hit all the points we want to hit?

OpenStudy (anonymous):

Hit the points

OpenStudy (amistre64):